# AB-Training: A Communication-Efficient Approach for Distributed Low-Rank Learning

### v1

@@ Abstract.
Communication bottlenecks hinder the scalability of distributed neural network training, particularly on distributed-memory computing clusters. To significantly reduce this communication overhead, we introduce AB-training, a novel data-parallel training method that decomposes weight matrices into low-rank representations and utilizes independent group-based training. This approach consistently reduces network traffic by 50% across multiple scaling scenarios, increasing the training potential on communication-constrained systems. Our method exhibits regularization effects at smaller scales, leading to improved generalization for models like VGG16, while achieving a remarkable 44.14 : 1 compression ratio during training on CIFAR-10 and maintaining competitive accuracy. Albeit promising, our experiments reveal that large batch effects remain a challenge even in low-rank training regimes.
Distributed training, Low-rank representation, Neural networks, Communication efficiency
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 53rd International Conference on Parallel Processing; August 12–15, 2024; Gotland, Swedenccs: Computing methodologies Neural networksccs: Computing methodologies Machine learning algorithmsccs: Networks Network performance analysisccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Batch learning
Figure 1. Average total interconnect traffic measured during the training of a Vision Transformer (ViT) B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$ with traditional, synchronous, full-rank data parallel training as implemented by PyTorch’s DistributedDataParallel (DDP) class and our AB training method. The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
[…]
Average total interconnect traffic measured during the training of a Vision Transformer B/16 on the ImageNet-2012 dataset for various node counts in $\mathrm{G}\mathrm{B}\mathrm{/}\mathrm{s}$. The global batch size for each training step is 4,096 images and is split evenly between the GPUs on each node.
1. Introduction
@@
The relentless pursuit of predictive performance in machine learning has led to neural networks of unprecedented scale and complexity. These massive state-of-the-art models excel in tasks ranging from image classification and natural language processing to scientific simulations. However, this performance comes at a steep computational cost, as training large models often demands large datasets. These datasets often exceed the capabilities of individual devices, necessitating distributed training approaches.
@@
Data parallelism (DP) is the most common distributed training technique for neural networks (Ben-Nun and Hoefler, 2019). It replicates the model across multiple devices, each of which processes a subset of the data, known as the local batch, before aggregating the gradients and updating model parameters. A persistent challenge in data parallel (DP) training is the communication bottleneck created by synchronizing large model representations across compute nodes (Niu et al., 2011). This bottleneck severely limits scalability, particularly on distributed-memory computing systems with potential network constraints. Large batch effects (Goyal et al., 2018), pose an additional challenge for maintaining predictive performance, as accuracy degrades once the global batch size crosses a critical threshold.
@@
To address these challenges, we propose a novel low-rank training method based on the singular value decomposition (SVD) and a hierarchical training scheme designed to significantly reduce communication overheads in distributed settings. Our method leverages the observation that low-rank representations can promote regularization (Coquelin et al., 2024) and minimize the amount of data transferred. Key contributions of our work include:
  • We find a 50% reduction in network traffic during distributed training on average as compared to traditional synchronous DP training, enabling experimentation with larger batch sizes and potentially accelerating training within distributed computing environments without high-speed interconnects.
@@
  • Our method demonstrates improved regularization effects at smaller scales, particularly on the Vision Transformer (ViT), leading to better predictive performance on unseen data. However, it also sheds light onto the challenge of reducing large batch effects without hyperparameter tuning.
@@
  • In an ideal scenario, AB training achieves a compression ratio of 44.14 : 1 for the VGG16 (Liu and Deng, 2015) model when trained on CIFAR-10 (Krizhevsky, 2009) while maintaining a competitive accuracy. In more standard scenarios, it achieved compression values ranging from 1.19 : 1 to 2.54 : 1 when exceeding the accuracy of traditional DP training.
@@
Our work highlights promising directions for how to fully harness the potential of low-rank training in massively scaled scenarios, with implications for both distributed-memory computing and the broader ML community.
2. Related Work
2.1. Distributed Training of Neural Networks
@@
The computational demands of training large neural networks have driven the development of distributed training paradigms, with data parallelism as a cornerstone approach. In a traditional data parallel (DP) training scheme, the model is replicated across multiple devices, each processing a subset of the training data independently. The data subset that one model instance processes is the local batch, while the global batch is the set of all local batches. After each forward-backward pass, gradients are aggregated across all model instances before the model is updated by the optimizer. As the number of compute resources and the model size increase, communication overheads incurred during gradient synchronization become a significant constraint, particularly in systems without high-bandwidth interconnects. Other parallelization strategies, such as model parallelism, introduce alternative ways to distribute training across devices, but come with their own communication and load-balancing complexities.
@@
Several techniques have been proposed to mitigate the communication bottleneck in data parallel training, including gradient accumulation, topology-aware communication patterns, asynchronous methods, and gradient compression (Vogels et al., 2019; Abrahamyan et al., 2022). Methods like H-SGD (Lin et al., 2020b) and DASO (Coquelin et al., 2022) utilize localized synchronization within smaller groups between global updates, potentially reducing communication costs by leveraging the network’s topology. Asynchronous approaches typically make use of a parameter server for aggregating gradients (Niu et al., 2011) and often require extensive hyperparameter tuning. Despite these advancements, the scalability of data parallelism remains constrained by communication bottlenecks and large batch effects, where increasing the global batch size negatively impacts generalization performance.
2.2. Low-Rank Neural Network Training
@@
A promising avenue for reducing both the memory footprint and computational cost of neural networks lies in low-rank training. Singular Value Decomposition (SVD) plays a crucial role in this domain. By factoring a weight matrix $W$ into three matrices, $U$, $\Sigma$, and $V^{T}$, and retaining only the $k$ largest singular values, a low-rank approximation with significantly fewer parameters can be obtained given a sufficiently large $k$ value. Note that aggressive compression can lead to a loss of approximation quality.
@@
Several methods leverage SVD for compressing neural networks during training. Some train directly on the $U$, $\Sigma$, and $V$ matrices (Schotthöfer et al., 2022; Coquelin et al., 2024), while others create different representations using the SVD and training on them (Liebenwein et al., 2021; Yang et al., 2020a). Both approaches can offer regularization-like benefits, potentially improving generalization (Winata et al., 2020; Cahyawijaya et al., 2021; Phan et al., 2020).
@@
Gradient compression using SVD has found application in federated learning scenarios (Wang et al., 2023b), where bandwidth constraints necessitate reduced communication. While these techniques address computational challenges in neural network training, their integration within distributed environments to specifically optimize communication remains an active area of research.
Low-rank factorizations can heavily reduce the communication required during distributed training. Existing approaches simply train low-rank models in parallel, but do not make any efforts to use the distributed system to their advantage (Coquelin et al., 2024; Wang et al., 2023a, 2021). To the best of our knowledge, there exist no methods that leverage both distributed architectures and low-rank representations to train neural networks.
3. Methodology
@@
Data parallel training of neural networks is constrained by the the communication bottleneck caused by synchronizing massive model representations across compute nodes and large batch effects, a phenomenon where accuracy degrades at large global batch sizes. In an attempt to tackle these challenges, our proposed training method integrates low-rank weight representations within a data parallel framework. It leverages the reduced network capacity stemming from topologically informed communication patterns and low-rank training while escaping local minima using group training.
We utilize SVD to decompose weight matrices into low-rank representations, significantly reducing the amount of data communicated during training. We decompose the weights of a given layer as
(1) $W_{m\times n}=A_{m\times k}B_{k\times n}=\left(U_{m\times k}\Sigma_{k\times k}^{1/2}\right)\left(\Sigma_{k\times k}^{1/2}V_{k\times n}^{\top}\right)$
where $U$, $\Sigma$, and $V$ are determined by the SVD of $W$ and $k<\texttt{min}\left(m,n\right)$. Using a low-rank representation in training has the added benefit of reducing overfitting, a common problem in training neural networks (Yang et al., 2020b; Hu et al., 2021).
@@
Large batch effects can be trickier to deal with. Consider the standard batch stochastic gradient descent (SGD) optimization step
(2) $W_{t}=W_{t-1}-\frac{\eta}{N}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{t-1}\right),$
@@
(3) $W_{k}=W_{0}-\frac{\eta}{B}\sum_{s=1}^{k-1}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{s}\right)$
With a larger batch size, $B$, the magnitude of the gradients shrinks. This removes the randomness inherent to SGD’s success and makes the network more prone to getting stuck in local minima (Chaudhari et al., 2019).
To mitigate this effect, we draw motivation from ensemble approaches. By dividing the DP model replicas into subgroups to be trained on independent data subsets, we essentially form an ensemble of models. Individually, these groups converge towards different local minima compared to full-system training, promoting exploration of the loss landscape.
After training in subgroups, we average the independently trained models. This merging strategy has seen some success in federated learning methods (Giuseppi et al., 2022). Compared to federated use-cases, which typically deal with models trained independently on biased datasets, AB training benefits from randomly sampling from the same dataset and thus has unbiased batches during our independent training phases.
During the independent training phases, each group trains on either $A$ or $B$, while the other component remains fixed. As the orthogonal component of $W$ is stable later in training (Coquelin et al., 2024), we only train on one of $A$ or $B$ in an attempt to reduce the divergence between the independent groups. At the end of the independent training phase, we average $A$ and $B$ across all workers to merge the independently trained models together.
@@
Distributed warmup:
Initial training with full-rank data parallelism allows the weights to adapt rapidly away from their random initialization. To avoid early divergence, we employ a traditional, full-rank data parallel warmup phase for a given number of iterations (warmupSteps).
Independent AB decomposition:
Each model instance independently computes an AB decomposition of its weight matrices as per Equation 1. This AB decomposition allows us to represent the weight matrices more compactly, reducing the amount of data to be communicated during training. The rank of this approximation is determined by a user-provided hyperparameter (sigmaCutoff) times the largest singular value. Values smaller than this are removed.
Group-only training:
Half of the independent groups will train the $A$ matrix, denoted as the $A$ groups, while the other groups will train $B$, the $B$ groups. For a given number of iterations (numABSteps), processes in the $A$ group only train the $A$ matrices, while processes in the $B$ group only train the $B$ matrices, the matrix not being trained remains frozen in this phase.
Synchronization and update:
We average both the trained and frozen $A$ and $B$ matrices independently to synchronize the model instances across all processes. This is crucial for ensuring that the overall model remains consistent across all compute nodes. Afterwards, we reconstruct the full-rank weight matrices from the updated $A$ and $B$ components.
Full-rank rebound:
We then train the full-rank network with traditional data parallelism for a short period, which we refer to as the fullRankReboundSteps. This phase promotes convergence and helps mitigate potential accuracy degradation observed in low-rank training scenarios.
A formalized algorithm is shown in Algorithm 1 and the method diagram is shown in Figure 2.
Figure 2. The AB training procedure.
[…]
The AB training procedure. See pseudocode or description in text.
@@ 5      $groupTrainTarget\leftarrow B$
11             $A,B\leftarrow\texttt{abDecomposition}(W)$triangleright Equation 1
@@ 21             $W\leftarrow AB$
Popular neural network optimizers like AdamW (L. and H., 2018) rely on second-order derivative approximations. When we change the network’s structure (e.g., between low-rank and full-rank representations), these optimizer states become incompatible due to their dimensionally. To address this, we use a “learning rate rebound” strategy. This involves reducing the learning rate to near-zero and gradually increasing it back to its scheduled value over a user-provided number of steps, allowing the optimizer to adjust to the new parameter representations. We do a learning rate rebound after each time the network’s parameters change shape.
4. Experimental Evaluation
To evaluate the effectiveness of our AB training approach, we conducted experiments on established neural network architectures. Our primary objectives were to demonstrate significant reductions in communication overhead, potential regularization benefits, scalability, and achievable compression during training. We focused on the ResNet-50 (He et al., 2016) and Vision Transformer (Dosovitskiy et al., 2021) B/16 models trained on the ImageNet-2012 dataset (Krizhevsky et al., 2012) for image classification as well as VGG16 (Liu and Deng, 2015) trained on CIFAR-10 (Krizhevsky, 2009). We compare our results against traditional DP training as implemented by PyTorch’s DDP package, referred to as ‘Traditional DDP’ or ‘Trad. DDP’, using the same hyperparameters (HPs). We also show results for the AB training setup without any individual trained groups which are denoted as ‘No Group’. In these experiments, the $B$ matrix is trained using traditional DP training in place of the full-rank weights. The ViT HPs are those used in the original source, while the ResNet-50 parameters utilize the same learning rate scheduler and the default values for the AdamW (L. and H., 2018) optimizer. The VGG16 HPs are as shown in (Coquelin et al., 2024).
We analyze our training method using two scaling strategies: constant local batch size scaling and constant global batch size scaling. In constant local batch size scaling, the global batch size increases with the number of GPUs. In constant global batch size scaling, we maintain a constant global batch size while proportionally decreasing the local batch size as we increase the number of GPUs. This emphasizes the communication overhead reductions achieved by our method. Measurements represent the average of three runs, each of which utilizes a different random seed. Models are initialized using orthogonal initialization (Hu et al., 2020).
4.1. Computational Environment

### v2

@@ Abstract
Communication bottlenecks severely hinder the scalability of distributed neural network training, particularly in high-performance computing (HPC) environments. We introduce AB-training, a novel data-parallel method that leverages low-rank representations and independent training groups to significantly reduce communication overhead. Our experiments demonstrate an average reduction in network traffic of approximately 70.31% across various scaling scenarios, increasing the training potential of communication-constrained systems and accelerating convergence at scale. AB-training also exhibits a pronounced regularization effect at smaller scales, leading to improved generalization while maintaining or even reducing training time. We achieve a remarkable 44.14 : 1 compression ratio on VGG16 trained on CIFAR-10 with minimal accuracy loss, and outperform traditional data parallel training by 1.55% on ResNet-50 trained on ImageNet-2012. While AB-training is promising, our findings also reveal that large batch effects persist even in low-rank regimes, underscoring the need for further research into optimized update mechanisms for massively distributed training.
Index Terms:
neural networks, data parallel training, low-rank decomposition, singular value decomposition
I Introduction
@@
The relentless pursuit of predictive performance in machine learning has led to neural networks of unprecedented scale and complexity. These massive state-of-the-art models excel in tasks ranging from image classification and natural language processing to scientific simulations. However, this performance comes at a steep computational cost, as training large models often demands large datasets. The sizes of modern models and datasets necessitates the usage of distributed training approaches.
@@
Data parallelism (DP) is the most common distributed training technique for neural networks [1]. It replicates the model across multiple devices, each of which processes a subset of the data, known as the local batch, before aggregating the gradients and updating model parameters. A persistent challenge in data parallel (DP) training is the communication bottleneck created by synchronizing large model representations across compute nodes [2].
As the accelerators of neural networks increase in computing power and on-board memory, models finish the forward-backward pass faster. In the standard implementations of DP training, the model gradients are eagerly synchronized once they are calculated during the backward pass. As models get larger and accelerators get faster, this synchronization happens more frequently and with more data. To solve this, one must either increase the available interconnect bandwidth or reduce the rate at which packets are sent to other nodes.
This bottleneck severely limits scalability, particularly on distributed-memory computing systems with potential network constraints. Furthermore, DP training increases the batch size used to calculate the gradients themselves which, at large scales, reduces the generalizablity of the network [3]. These large batch effects pose an additional challenge for maintaining predictive performance, as accuracy degrades once the global batch size crosses a critical threshold.
@@
To address these challenges, we propose a novel low-rank training method based on the singular value decomposition (SVD) and hierarchical training specifically designed to reduce the interconnect traffic in distributed settings. Our method leverages the fact that low-rank representations can reduce the amount of data transferred [4], as well as the observation that low-rank representations can promote regularization [5]. Key contributions of our work include:
  • Through the usage of independent training groups and low-rank representations, we reduce the network traffic during distributed training by an average of70.31% as compared to traditional synchronous DP training.
@@
  • Our method demonstrates improved regularization effects at smaller scales, particularly on the Vision Transformer (ViT), leading to better predictive performance on unseen data.
@@
  • In an ideal scenario, AB training achieves a compression ratio of 44.14 : 1 for the VGG16 [6] model when trained on CIFAR-10 [7] while maintaining a competitive accuracy. In more standard scenarios, it achieved compression values ranging from 1.19 : 1 to 2.54 : 1 when exceeding the accuracy of traditional DP training.
@@
Our work highlights promising directions for how to harness the potential of low-rank training in large-scale scenarios, with implications for both distributed-memory computing and the broader ML community. We provide an open-source implementation of our method, as well as the configurations to reproduce the experiments shown [^1] Will be made available upon publication.
II Related Work
II-A Distributed Training of Neural Networks
@@
The computational demands of training large neural networks have driven the development of distributed training paradigms, with data parallelism as a cornerstone approach. In a traditional data parallel (DP) training scheme, the model is replicated across multiple devices, each processing a unique subset of the training data independently. The data subset that one model instance processes is the local batch, while the global batch is the set of all local batches. After each forward-backward pass, gradients are aggregated across all model instances before the model is updated by the optimizer.
As the number of compute resources and the model size increase, communication overheads incurred during gradient synchronization become a significant constraint. Other parallelization strategies, such as model parallelism, introduce alternative ways to distribute training across devices, but come with their own communication and load-balancing challenges.
@@
Several techniques have been proposed to mitigate the communication bottleneck in data parallel training, including gradient accumulation, topology-aware communication patterns, asynchronous methods, and gradient compression [4, 8]. Methods like H-SGD [9] and DASO [10] utilize localized synchronization within smaller groups between global updates, reducing communication costs by leveraging the system’s network topology. Asynchronous approaches typically make use of a parameter server for aggregating gradients [2] and often require extensive hyperparameter tuning [1]. Despite these advancements, the scalability of data parallelism remains constrained by communication bottlenecks and large batch effects, where increasing the global batch size negatively impacts generalization performance.
II-B Low-Rank Neural Network Training
@@
A promising avenue for reducing both the memory footprint and computational cost of neural networks lies in low-rank training. Singular Value Decomposition (SVD) plays a crucial role in this domain. By factoring a neural network’s weight matrix $W$ into three matrices (an orthogonal basis $U$, a diagonal matrix of singular values $\Sigma$, and an orthogonal basis $V^{T}$) and retaining only the $k$ largest singular values, a low-rank approximation with significantly fewer parameters can be obtained given a sufficiently large $k$. However, aggressive compression can lead to a loss of approximation quality.
@@
Several methods leverage SVD for compressing neural networks during training. Some train directly on the $U$, $\Sigma$, and $V$ matrices [11, 5], while others create different representations using the SVD and training on them [12, 13]. Both approaches can offer regularization-like benefits, potentially improving generalization [14, 15, 16]. LoRA [17], and methods similar to and inspired by it, utilize low-rank representations alongside the frozen full-rank weights for fine-tuning.
@@
Gradient compression using SVD has found application in federated learning scenarios [18], where bandwidth constraints necessitate reduced communication. While these techniques address computational challenges in neural network training, their integration within distributed environments to specifically optimize communication remains an active area of research. More precisely, existing approaches simply train low-rank models in parallel, but do not make any efforts to use the distributed system to their advantage [5, 19, 20].
II-C Pruning
Pruning aims to reduce model size by removing irrelevant parameters. It is the primary competition to low-rank methods for reducing the size of neural networks. One well-known method is Iterative Magnitude Pruning (IMP), which involves repeatedly training a network, removing small-magnitude weights, and retraining from the initial state[21]. While this process can result in highly compressed models., unstructured pruning methods like IMP often lead to irregular sparsity patterns, making it difficult to leverage specialized hardware optimized for sparse computation. Structured pruning techniques target specific components of the network. For example, in convolutional layers removing filters [22, 23, 24] or channels [25, 26, 27] can maintain traditional dense parameter layouts to make use of efficient, pre-existing acceleration methods.
III Methodology
@@
Data parallel training of neural networks is constrained by the the communication bottleneck caused by synchronizing massive model representations across compute nodes and large batch effects, a phenomenon where accuracy degrades at large global batch sizes. In an attempt to tackle these challenges, our proposed training method integrates low-rank weight representations within a hierarchical data parallel framework. It leverages the reduced network capacity stemming from topologically informed communication patterns and low-rank training while also escaping local minima using group training.
@@
Working with large batches can have unintended consequences. Consider the standard batch stochastic gradient descent (SGD) optimization step
$W_{t}=W_{t-1}-\frac{\eta}{B}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{t-1}\right),$ (1)
@@
$W_{k}=W_{0}-\frac{\eta}{B}\sum_{s=1}^{k-1}\sum_{i=1}^{B}\nabla Q_{i}\left(W_{s}\right)$ (2)
With a larger batch size, $B$, the magnitude of the gradients shrinks. This removes the randomness inherent to SGD’s success and makes the network more prone to getting stuck in local minima [28, 29].
To mitigate this effect, we draw motivation from the fact that the noise inherent to the gradients in minibatch SGD is an important factor in why it generalizes better than large batch gradient descent [29]. During training we divide the workers into small sub-groups; each of which train the model independently for a number of steps. By dividing the DP model replicas into subgroups to be trained on independent data subsets, we effectively form an ensemble of models. Individually, these groups benefit from the improved generalization of minibatch SGD by using the noisier gradients to explore of the loss landscape to find minima which would otherwise be missed by large-batch SGD. Additionally, training in independent groups heavily reduces the network traffic.
To further reduce the amount of data communicated during training, we utilize SVD to decompose weight matrices into low-rank representations. We represent the weights of a given layer as
$W_{m\times n}=A_{m\times k}B_{k\times n}=\left(U_{m\times k}\Sigma_{k\times k}^{1/2}\right)\left(\Sigma_{k\times k}^{1/2}V_{k\times n}^{\top}\right)$ (3)
where $U$, $\Sigma$, and $V$ are determined by the SVD of $W$ and $k<\texttt{min}\left(m,n\right)$. In the case that the weight matrix has more than two dimensions, the trailing dimensions are flattened to form a two-dimensional representation to be used as $W$. If the shape of the matrix $W$ is unfavorable ($n>m$) we transpose $W$ before the decomposition.
One of the most obvious methods for training a low-rank network would be to train directly on the $\Sigma$ matrix. However, it is not beneficial to have any trainable weights set to zero during training as these connections are effectively dead during backpropagation. Since the off-diagonal values of $\Sigma$ (as dictated by SVD) are zero, training on only $\Sigma$ will have the effect of freezing the orthogonal vectors of $U$ and $V$. To allow the network to modify the orthogonal vectors and to maintain a dense representation, we decompose the network as shown in Equation 3. Interestingly, training a network using low-rank representations has the added benefit of reducing overfitting, a common problem where the performance on the training dataset is much higher than on the validation or test datasets [13, 30].
During the independent training phase, half of the sub-groups train $A$ while the others train $B$; in this phase the other component remains fixed. This ensures that at least a portion of the orthogonal vectors inherent to $W$ are fixed in time, reducing the chance that the groups greatly diverge and increases the gradient noise due to the smaller batch size. With this method, the groups can more easily learn to modify the full low-rank representations of the network’s parameters.
At the end of the independent training phase, we average $A$ and $B$ across all workers to merge the independently trained models. This merging strategy has seen success in federated learning [31], a use-case which typically deals with models trained independently on biased datasets. As we use unbiased batches during the independent training phase, we expect the negative aspects of this average to be minimal.
@@
  • Full Rank DP Warmup: Initial training with full-rank data parallelism allows the weights to move rapidly away from their random initialization. To avoid early divergence, we employ a traditional, full-rank data parallel warmup phase for a given number of iterations (warmupSteps).
  • Independent AB decomposition: Each model instance independently computes an AB decomposition of its weight matrices as per Equation 3. This AB decomposition allows us to represent the weight matrices more compactly, reducing the amount of data to be communicated during training. The rank of this approximation is determined by a user-provided hyperparameter (sigmaCutoff) and the largest singular value. Values smaller than the largest singular value times sigmaCutoff are removed.
  • Group training: Half of the independent groups will train the $A$ matrix, denoted as the $A$ groups, while the other groups will train $B$, the $B$ groups. For a given number of iterations (numABSteps), the $A$ group only trains the $A$ matrices and the $B$ group only trains the $B$ matrices, the matrices not being trained are frozen.
  • Synchronization and update: We average both the trained and frozen $A$ and $B$ matrices to synchronize the model instances across all processes. Afterwards, we reconstruct the full-rank weight matrices from the updated $A$ and $B$ components.
  • Full-rank rebound: We then train the full-rank network with traditional data parallelism for a number of iterations (fullRankReboundSteps). This phase promotes convergence and helps mitigate potential accuracy degradation observed in low-rank training scenarios.
A formalized algorithm is shown in Algorithm 1 and a method diagram is shown in Figure 1.
Figure 1: A UML diagram of the AB training procedure.
@@ 5      $groupTrainTarget\leftarrow B$
11             $A,B\leftarrow\texttt{abDecomposition}(W)$triangleright Eq. 3
@@ 21             $W\leftarrow AB$
Popular neural network optimizers like AdamW [32] rely on second-order derivative approximations. When we change the network’s structure, e.g., between low-rank and full-rank representations, the optimizer states used in these approximations become incompatible due to their dimensionally. To address this, we use a learning rate rebound strategy. This involves reducing the learning rate to near-zero and gradually increasing it back to its scheduled value over a user-provided number of steps, allowing the optimizer to adjust to the new parameter representations. We do a learning rate rebound whenever the network’s parameters change shape.
IV Experimental Evaluation
To evaluate the effectiveness of our AB training approach, we conducted experiments on well-studied neural network architectures. Our primary objectives were to demonstrate significant reductions in interconnect traffic, potential regularization benefits, scalability, and achievable compression during training. We focused on the ResNet-50 [33] and Vision Transformer [34] B/16 models trained on the ImageNet-2012 dataset [35] for image classification as well as VGG16 [6] trained on CIFAR-10 [7]. We compare our results against traditional full-rank DP training as implemented by PyTorch’s DistributedDataParallel (DDP) class, referred to as ‘Traditional DDP’ or ‘Trad. DDP’, using the same hyperparameters (HPs). In these experiments, the $B$ matrix is trained during the group phase using traditional DP training. The ViT HPs are those used in the original source, while the ResNet-50 parameters utilize the same learning rate scheduler as the ViT and the default values for the AdamW [32] optimizer. The VGG16 HPs are as shown in [5].
We analyze our training method using two scaling strategies: constant local batch size scaling and constant global batch size scaling. In constant local batch size scaling, the global batch size increases with the number of GPUs. In constant global batch size scaling, we maintain a constant global batch size while proportionally decreasing the local batch size as we increase the number of GPUs. This emphasizes the interconnect traffic reduction achieved by our method. To disentangle the effects of training with independent groups and those of low-rank training, we show measurements for AB training with and without groups. For the no-group measurements, the workers train the $B$ matrix with traditional DP methods and denoted as ‘AB - No Groups’. At a fixed scale, we compare AB training with and without groups to other pruning and low-rank methods. Measurements represent the average of three runs, each of which utilizes a different random seed. Models are initialized using orthogonal initialization [36].
IV-A Computational Environment